### v1

# Free3D: Consistent Novel View Synthesis without 3D Representation

Abstract
We introduce Free3D, a simple approach designed for open-set novel view synthesis (NVS) from a single image. Similar to Zero-1-to-3, we start from a pre-trained 2D image generator for generalization, and fine-tune it for NVS. Compared to recent and concurrent works, we obtain significant improvements without resorting to an explicit 3D representation, which is slow and memory-consuming or training an additional 3D network. We do so by encoding better the target camera pose via a new per-pixel ray conditioning normalization (RCN) layer. The latter injects pose information in the underlying 2D image generator by telling each pixel its specific viewing direction. We also improve multi-view consistency via a light-weight multi-view attention layer and multi-view noise sharing. We train Free3D on the Objaverse dataset and demonstrate excellent generalization to various new categories in several new datasets, including OminiObject3D and GSO. We hope our simple and effective approach will serve as a solid baseline and help future research in NVS with more accuracy pose. The project page is available at https://chuanxiaz.com/free3d/.
Objaverse [18]
OmniObject3D [81]
GSO [20]
Input ViewGenerated Views by our Free3D Input ViewGenerated Views by our Free3D
Figure 1: Free3D novel view synthesis in an open-set setting. Given a single input view, our approach synthesizes consistent $360^{\circ}$ videos accurately and with no explicit 3D representation. Trained on Objaverse only, it generalizes well to new datasets and categories.
1 Introduction
Novel view synthesis (NVS) has developed rapidly in the recent past [5, 49, 6, 11, 38], fuelled in part by NeRF [48]. However, many such NVS methods require to optimize a new model from scratch for each scene, and are thus impractical in many applications. Inspired by generative models [39, 25, 67], other NVS methods generalize beyond a single scene to a specific object category, or even to unstructured datasets of objects [50, 34, 85, 32, 76, 47, 69]. Remarkably, these methods, by utilizing the data prior, can generate new views starting from a single image. However, they trade off some quality in order to achieve generality, which is still insufficient to build a 3D foundation model. Going beyond, in this work we consider the NVS problem in an open-set setting, generalizing beyond a single object, category, and dataset, yet still utilizing a single image as input, and aiming at obtaining high-quality outputs.
There are two primary quality targets: (i) The output must accurately reflect the pose of the target cameras, and (ii), when several views of the same object are generated, they must be multi-view consistent. In order to achieve high 3D accuracy and consistency, recent and concurrent methods [42, 44, 63, 83] use a 3D representation of the reconstructed object, usually on top of a pre-trained 2D generative model [57]. It works well but adds complexity.
In this paper, we introduce Free3D, a simple, flexible, and efficient method that can also achieve consistent NVS results without the need to rely on a resource-intensive 3D representation. Zero-1-to-3 [43] is perhaps the best-known example of such a 3D-free NVS system. Like it, Free3D builds upon a pre-trained 2D generative model like Stable Diffusion [57], trained on a large-scale dataset (over 5 billion images [62]), as a data prior. The prior knowledge contained in such a 2D generator is extremely important to be able to ‘guess’ plausible novel views of arbitrary new objects, which is inherently highly ambiguous. However, we show empirically that Zero-1-to-3, has, in practice, poor camera pose control, and, when tasked with generating multiple views, not very consistent. The latter is unavoidable in their design because each view is sampled independently from scratch and, owing to the ambiguity of reconstruction, there is no reason why compatible images would be generated each time. We refer readers to the supplemental videos for a comprehensive comparison.
To mitigate these issues, we first show that much better camera control can be achieved by switching to a different representation of camera pose. Specifically, we introduce a ray conditioning normalization (RCN) (Fig. 2 (b)), which tells each pixel its viewing direction. This is a distributed representation, which should be contrasted to the concentrated camera representation used in [43]. They pass the camera elevation, azimuth, and distance as language-like tokens that may be difficult for the network to interpret and utilize [61]. In contrast, with our RCN layer, we show how to effectively incorporate this per-pixel ray information in an existing text-to-image diffusion model, which empirically leads to significantly more accurate NVS in our experiments (Tabs. 1 and 2). It is also reminiscent of the design of methods like NeRF [48] and LFNs [66], which also work by processing individual rays.
While RCN leads to more accurate camera control, it cannot improve multi-view consistency by itself. To improve the latter, we introduce an additional pseudo-3D cross-view attention module (Fig. 2(c)), inspired by video diffusion models [64, 30, 8, 26, 80], that fuses information across all views instead of processing each view independently. Furthermore, we use multi-view noise sharing when generating the different views of the object, which further enhances consistency due to the continuity of the denoising function, reducing aleatoric variations from view to view.
We benchmark Free3D against recent and concurrent state-of-the-art methods [43, 44, 77, 17] on open-set NVS. Although the model is trained on only one dataset, it generalizes well to all recent NVS benchmark datasets, including Objaverse [18], OmniObject3D [81], and Google Scanned Object (GSO[20]. A thorough experimental assessment shows that our approach consistently outperforms existing methods, both quantitatively and qualitatively.
To summarise, with Free3D we make the following contributions: (i) We introduce the ray conditioning normalization (RCN) layer and show that representing the camera by utilizing a combination of distributed ray conditioning and concentrated pose tokens significantly improves pose accuracy in NVS. (ii) We show that a small multi-view attention module can improve multi-view consistency by exchanging information between views at a low cost. (iii) We find that multi-view noise sharing between different views further improves consistency. (iv) We demonstrate empirically that Free3D achieves consistent NVS without needing a 3D representation and outperforms the existing state-of-the-art models on both pose accuracy and view consistency.
2 Related Work
Per-Scene NVS.
Early NVS works relied on epipolar geometry to interpolate between different views of the same scene [14, 16]. A recent breakthrough was to represent 3D scenes as implicit neural fields, as proposed by SRN [65], DeepSDF [51], NeRF [48] and LFN [66], and further improved in follow-ups [87, 5, 6, 24, 11, 49, 38]. Even so, data efficiency, generalizability, and robustness remain a limitation: such systems require multiple views to learn the 3D representation from scratch of every single scene.
To bypass the need for multiple input views, DreamFusion [52] proposed to distill 3D models from a large-scale pre-trained 2D diffusion model [58]. RealFusion [46] extended the latter to single-view image reconstruction by adding the input image as a constraint during distillation. Several follow-up works [70, 54, 61, 53] achieved further improvements to the resolution and quality of the resulting 3D assets. Although these models are open-ended [46], they require lengthy per-scene optimisation.
Refer to caption(a) Muti-View Diffusion Pipeline(b) Ray Conditional Normalisation(c) Pseudo 3D Cross-Attention
Figure 2: The overall pipeline of our Free3D. (a) Given a single source input image, the proposed architecture jointly predicts multiple target views, instead of processing them independently. To achieve a consistent novel view synthesis without the need for 3D representation, (b) we first propose a novel ray conditional normalization (RCN) layer, which uses a per-pixel oriented camera ray to module the latent features, enabling the model’s ability to capture more precise viewpoints. (c) A memory-friendly pseudo-3D cross-attention module is introduced to efficiently bridge information across multiple generated views. Note that, here the similarity score is only calculated across multiple views in temporal instead of spatial, resulting in a minimal computational and memory cost.
Category-Specific NVS.
Inspired by the success of 2D generation [39, 25, 67], some work [71, 50, 78, 15] built the autoencoder architecture for NVS. Driven by the effectiveness of light field rendering (LFR) [2, 1], other generalizable NVS approaches [68, 60, 59] query a network for colour of different rays. 3DiM [76] then introduced diffusion models into NVS. This is followed by RenderDiffusion [3], HoloDiffusion [36], SparseFusion [90], GeNVS [9], Viewset Diffusion [69] and LFD [82]. However, these methods train their models from scratch using pure 3D data. Because such datasets are not sufficient for open-set generalization, they are limited to one or few categories. Here we start from an open-ended 2D image generator instead.
Open-set NVS.
In order to operate in an open-set setting, Zero-1-to-3 [43] built on a 2D image generator trained on a large-scale image dataset [62] and fine-tune it on the Objaverse dataset [18]. While it has good generalizability, Zero-1-to-3 fails to achieve high pose accuracy, and their reconstructions are not very consistent across views. To mitigate these issues, other concurrent approaches either integrate 3D representations into the network [83, 44, 35] or train another auxiliary 3D network [42, 77, 63, 41]. These methods can output high-quality target views, but they are computationally expensive and use significant memory for training. More importantly, these methods do not address the issue of pose representation, which we find to be a key bottleneck in NVS. In contrast, our method is 3D-free and achieves comparable or better NVS quality, due to ray conditioning normalisation, multi-view attention, and multi-view noise sharing. A more recent concurrent work is Stable Video Diffusion [7], which also synthesizes multi-view videos. However, they are trained with fixed frames with limited elevation angles. In contrast, our Free3D is trained to handle arbitrary camera viewpoints, and ideally generates ${}^{\circ}$ spectrum, with flexible-frames for any elevation angles.
3 Method
Our goal is to learn a model $\Phi$ that, given as input a source image $x^{\text{src}}$ and a sequence of camera poses $\mathcal{P}=\{\mathbf{P}^{i}\}^{N}_{i=1}$, $\mathbf{P}^{i}=(\mathbf{K}^{i},\mathbf{R}^{i},\mathbf{T}^{i})$ comprising intrinsics $\mathbf{K}^{i}$, rotation $\mathbf{R}^{i}$ and translation $\mathbf{T}^{i}$, synthesizes corresponding novel views $\{x^{i}\}_{i=1}^{N}$ which are accurate and consistent without relying on an explicit 3D representation. We approach this problem by generating all the views together, conditioned on the source image $x^{\text{src}}$ and a series of camera parameters. In this way, the network has a chance to produce several consistent views together. The broad motivation here is that the 1D sequential representation has been used extensively in the 2D generation [74, 73, 75, 56, 13, 23, 89]. Analogously, it should be able to produce visually consistent video representing the underlying 3D structure using 2D sequential representation, i.e. multi-frames [7, 4]. However, we tackle it in a bidirectional manner [40, 22, 10, 89, 86], instead of sequentially auto-regressive generation.
To realize this goal, we must tackle two challenges: (i) ensuring that the model accurately captures the target view and (ii) ensuring that different views of the object are consistent in geometry and appearance. To achieve these, our framework, illustrated in Fig. 2, extends a 2D generator by injecting at each layer a ray-conditional normalization (RCN) layer (Sec. 3.1) as well as pseudo-3D attention layer (Sec. 3.2). It also utilizes multi-view noise sharing (Sec. 3.2). The former captures pose more accurately, whereas the last two improve consistency. Although several prior NVS works have started from 2D generators and also focused on improving multi-view consistency, with RCN we are the first (to the best of our knowledge) to use ray conditioning in this setting and obtain high pose accuracy without a 3D representation or an additional 3D network.
3.1 Ray Conditioning Normalization (RCN)
Ray conditioning was originally proposed for NVS by [66]. However, it has not been used for NVS models that build on 2D generators like Zero-1-to-3 [43] — in practice, doing so is essential for generalization due to the lack of very large-scale 3D data. Here, we thus develop a method to extend a pretrained 2D generator with ray conditioning.
Ray Conditioning Embedding.
Given a target view $\mathbf{P}=(\mathbf{K},\mathbf{R},\mathbf{T})$, for each pixel $(u,v)$ in the image, we define the Plücker coordinates $\bm{r}_{uv}=\phi(\bm{o},\bm{d}_{uv})=(\bm{o}\times\bm{d}_{uv},\bm{d}_{uv})\in\mathbb{R}^{6}$, of the ray going from the camera center $\bm{o}\in\mathbb{R}^{3}$ through the pixel, while $\bm{d}_{uv}=\mathbf{R}^{\top}(\mathbf{K}^{-1}(u,v,1)^{\top}-\mathbf{T})\in\mathbb{R}^{3}$ is the ray direction. This encoding was originally introduced by LFN [66]. It is invariant to shifting the camera along the ray, meaning that $\phi(\bm{o}+\lambda\bm{d},\bm{d})=((\bm{o}+\lambda\bm{d})\times\bm{d}_{uv},\bm{d}_{uv})=(\bm{o}\times\bm{d}_{uv},\bm{d}_{uv})=\phi(\bm{o},\bm{d})$, which matches the fact that light propagates in straight lines.
Ray Conditioning Architectures.
Our goal is to modify the denoising neural network $\hat{\epsilon}(z_{,}t,y)$ so that the conditioning information $y$ includes ray conditioning. We experiment with a number of different architectures to do so, proposing various ray conditioning layers:
  • Concatenation. Following Zero-1-to-3 [43], a natural choice is to concatenate the noised target $z_{t}^{\text{tgt}}$, original source embedding $z^{\text{src}}$, and ray conditioning at the input. Then, the input is now $(z_{t}^{\text{tgt}},z^{\text{src}},\bm{r})$ instead of $z_{t}^{\text{tgt}}$ alone.
  • Multi-scales concatenation. We further consider concatenating ray embeddings $\bm{r}$ to each intermediate layer in the UNet $\epsilon_{\theta}$. Note that each layer operates at a different resolution, so this amounts to injecting the information $\bm{r}$ at different scales.
  • Ray conditioning normalization (RCN). Finally, we propose to combine the adaptive layer norm [21, 31, 37, 89] with ray conditioning to modulate the image latents. Specifically, the activation latent $F_{i}$ of the $i$-th layer in the UNet $\epsilon_{\theta}$ is modulated by:
    $\operatorname{ModLN}_{\bm{r}}(F_{i})=\operatorname{LN}(F_{i})\cdot(1+\gamma)+\beta,$ (1)
    where $(\gamma,\beta)=\operatorname{MLP}^{\text{mod}}(\bm{r})$ are the learned scale and shift from ray conditioning $\bm{r}$. Such modulation is applied to each sub-module of UNet $\epsilon_{\theta}$ (Fig. 2 (a) & (b)).
Interestingly, while RCN works best, we show empirically (Tab. 3) that all such architectures lead to strong improvements. This confirms the importance of ray conditioning for the significant improvements, with the novel RCN emerging as the frontrunner in this progression.
Discussion.
RCN differs significantly from Zero-1-to-3 [43] and follow-ups [42, 44, 83, 77, 33] which coded $\mathbf{P}$ as global tokens. Ray-Cond [12] and LFD [82] also applied ray conditioning for NVS in GAN and Diffusion, respectively, but just concatenated it as additional channels. Besides, they are only trained for category-specific NVS. Similar to ray-casting [45, 48], RCN considers one ray per pixel. However, it eschews the computation of hundreds of samples per ray. Because of this, Free3D dramatically reduces the rendering time and memory consumption compared to the concurrent works of [44, 42, 83].
3.2 View Consistent Rendering
Given a source image $x^{\text{src}}$, our goal is to render a series of consistent novel views $x^{i}$. While the camera encoding technique of Sec. 3.1 significantly enhances the pose accuracy of our model, if images are sampled independently, they will almost never be visually consistent due to the intrinsic ambiguity of the reconstruction task. This is visible as temporal flickering in the rendered 3D video (as shown in supplemental videos). To remove or at least greatly mitigate this problem, we propose to sample images jointly as well as to share noise among them.
Multi-view attention.
We first adapt the frame attention, a well-established method in video diffusion models [64, 30, 8, 26, 80], to capture temporal dependencies across different views. As shown in Fig. 2(c), given a 5D latent $z\in\mathbb{R}^{B\times v\times c\times h\times w}$, we initially reshape it to $z\in\mathbb{R}^{(B\times h\times w)\times v\times c}$, resulting in $batch\times height\times width$ sequences at the length of $views$. Subsequently, this reshaped latent is passed through the pseudo-3D attention module to calculate the similarity across different views. Since this attention layer operates across views but separately for each spatial location, the computational and memory costs are quite low (Tab. 1). Similar to ray conditioning, we inject multi-view attention at each level of the UNet $\hat{\epsilon}_{\theta}$.
Multi-view noise sharing.
At test time, views are sampled using the backward process Eq. A.2 starting from a normally-distributed, yet randomly-sampled noise vector $x_{T}$. The idea is that the noise distribution is a representation of the data distribution, and models aleatoric uncertainty. In our case, while we are interested in generating multiple and diverse solutions, we still wish the different views to be internally consistent for each $360^{\circ}$ rendered video.
In order to reduce the variance between different views, we propose to start sampling each view from the same noise vector $x_{T}$. The network $\epsilon_{\theta}$ still generates different views because it is conditioned on the camera parameters. It can also generate different reconstructions by resampling $x_{T}$. However, the variance between views is thus reduced. Noise sharing can be justified by noting that the network $\hat{\epsilon}_{\theta}(z_{t},t,y)$ is a continuous function of both $z_{t}$ and $y$ [79]. Besides, different from [72], we do not specially set the noise schedule in different time steps $t$. More complex designs and training strategies have the potential to improve performance but are not the focus of this work.
3.3 Learning formulation
The learning objective is given by
$\mathcal{L}=\mathbb{E}_{(Z_{0},z^{\text{src}},\mathbf{P}),\epsilon,t}\left[\lVert\epsilon-\epsilon_{\theta}(Z_{t},t,y)\rVert^{2}_{2}\right],$ (2)
where each training sample $(Z_{0},z^{\text{src}},\mathbf{P})$ consist of $N$ encoded target views $Z_{0}=\{\mathcal{E}(x^{i})\}_{i=1}^{N}$, the encoded source view $z^{\text{src}}=\mathcal{E}(x^{\text{src}})$ and the viewpoints $\mathbf{P}$. The network is conditioned on the source view and cameras, i.e., $y=(z^{\text{src}},\mathbf{P})$ and, as in Fig. 2, estimates the noise for all target views jointly. This information is incorporated as discussed above via ray conditioning modulation; following [43], we also concatenate the input image code $z^{\text{src}}$ to $z_{t}$ along the channel dimension and add the CLIP [55] encoding.
Figure 3: Perceptual Path Length Consistency (PPLC). To partly compensate for the viewpoint change, the second image is rectified w.r.t. the first before comparison. To illustrate the importance of using rectification, the figure shows two objects in a large azimuth $\phi:57.6^{\circ}$. The top row shows to the left an ideally-rendered image pair, which however attains a large LPIPS loss due to the viewpoint change. To the right, rectification reduces this score. The bottom row shows the opposite, where a pair of incorrectly rendered views has its LPIPS loss increased by rectification.
3.4 Perceptual Path Length Consistency
To quantify the consistency across different novel views, recent methods [76, 43, 77, 83] trained NeRF models from the sampled views and evaluated them on the remaining views. However, this requires training a NeRF for each test instance, which is unfeasible for large-scale testing required for proper open-set NVS evaluation. Here, we propose to use instead the pairwise perceptual distance [88] to measure the consistency between generated views. In particular, we first subdivide the ${}^{\circ}$ rendering path into linear segments and then calculate the LPIPS score between two neighboring generated images $x^{i}$ and $x^{i+1}$. Naturally, these images differ due to the different viewpoints (Fig. 3 to row). We partly compensate for the viewpoint change by rectifying [27] the second image w.r.t. the first. We then define Perceptual Path Length Consistency (PPLC) of one rendered sequence $\{x^{i}\}_{i=1}^{N}$ as follows:
$l_{\text{pplc}}=\mathbb{E}\left[\frac{1}{\phi^{2}}\lVert\mathcal{F}(\text{Rect}(x^{i}))-\mathcal{F}(\text{Rect}(x^{i+1}))\rVert_{2}^{2}\right],$ (3)
where $\phi$ is the degree between views $x^{i}$ and $x^{i+1}$, which is set as $\phi=2\pi/50$, with the azimuth $7.2^{\circ}$ in all our video rendering. $\mathcal{F}$ is a pre-trained network to ensure the metric matches with human perceptual similarity judgment.
4 Experiments
Refer to caption(a) Input View(b) Target View(c) Zero1-to-3 [43](d) Zero123-XL [17](e) SyncDreamer [44](f) Consistent123 [77](g) Ours Free3D
Figure 4: Qualitative comparisons on Objaverse. Given a target pose, our Free3D significantly improves the accuracy of the generated pose compared to existing state-of-the-art methods. Note that Zero123-XL [17] is trained on the much larger Objaverse-XL dataset [17], which contains 10 million 3D objects. More comparisons are provided in the supplement Figs. C.1 and C.2.
Method Objaverse[18]
[…]
Zero-1-to-3[43] 0.8462 0.0938 1.52 18.84 3s/44s
Zero123-XL[17] 0.8339 0.1098 1.67 25.61 3s/44s
SyncDreamer[44] 0.8063 0.1910 7.57 16.32 25s/77s
Consistent123[77] 0.8530 0.0913 1.48 17.89 4s/63s
Ours (Free3D) 0.8620 0.0784 1.21 10.82 3s/52s
Table 1: Comparison with SoTA methods on all 7,729 objects in Objaverse test-set. Recent works like [76, 43, 17, 77] were originally evaluated using a subset of this data only due to the cost of training additional 3D models. Unlike them, we directly evaluate the model on whole test-set without using additional 3D networks. The inference time is for rendering a single target view and a $360^{\circ}$ video, respectively.
We compare Free3D to state-of-the-art open-ended single-image NVS methods and also ablate our design choices.
4.1 Experimental Details
Datasets.
For fairness, our model is trained using the exact same protocol as Zero-1-to-3 [43]. They render multiple views for 772,870 objects from Objaverse dataset [18]. We use the identical test split as they do. To assess how well our mode generalizes to other datasets, and how it compares to other models, we consider two more datasets: OmniObject3D [81] and Google Scanned Objects (GSO[20], which contain real-life scanned objects. Since we do not use these datasets for training at all, we use the entirety of OmniObject3D and GSO objects for evaluation (6,000 and 1,030 objects, respectively).
Metrics.
We follow [43, 77, 44] and assess the NVS quality by comparing the generated images and the ground-truth views at different levels of granularity, including pixel-level Peak Signal-to-Noise Ration (PSNR), patch-level Structure SIMilarity index (SSIM) and feature-level Learned Perceptual Image Patch Similarity (LPIPS) [88] [^1] https://github.com/richzhang/PerceptualSimilaritysqueeze”-net., and dataset-level Fréchet Inception Distance (FID) [28] [^2] https://github.com/GaParmar/clean-fidclip_vit_b_32”-net.. Some of these metrics look at the statistics of individually-reconstructed images rather than exact reconstructions, which is the correct approach given that the reconstruction task is inherently ambiguous; however, for the same reason, they are not suitable for assessing multi-view consistency. Instead, we use PPLC (Sec. 3.4) to evaluate consistency. To do so, we generate 50-frame videos along a pre-defined circular camera trajectory for each object and then compute the PPLC score between neighbouring frames.
Baselines.
We compare Free3D to the state-of-the-art Zero-1-to-3 [43]${}_{\text{ICCV'2023}}$ and the follow-ups Zero123-XL [17]${}_{\text{NeurIPS'2023}}$, SyncDreamer [44]${}_{\text{arXiv'2023}}$ and Consistent123 [77]${}_{\text{arXiv'2023}}$, which are built on Zero-1-to-3. While noticed other concurrent works in arXiv for NVS [84, 35, 83], but, as no codes are available yet, we could not re-implement and re-train all of them for comparison.
OmniObject3D [81]
GSO [20]
(a) Input View(b) Target View(c) Zero-1-to-3[43](d) Zero123-XL[17](e) SyncDreamer[44](f) Consistent123[77](g) Ours Free3D
Figure 5: Qualitative comparisons on OmniObject3D (top two rows) and GSO (bottom two rows) dataset. Interestingly, exciting methods cannot deal with unconventional objects, such as the “pie” in the first row, while our Free3D is still robust for such a challenging scenario. More comparisons are provided in supplemental Figs. C.3 and C.4.
Method OmniObject3D [81] GSO [20]
[…]
Zero-1-to-3 [43] 16.84 0.7813 0.1321 1.73 24.58 19.65 0.8501 0.0758 3.24 33.15
Zero123-XL [17] 17.11 0.7818 0.1291 1.51 21.33 20.43 0.8589 0.0706 3.23 28.03
SyncDreamer [44] 17.00 0.7941 0.1442 6.58 11.49 14.72 0.7835 0.1533 8.65 9.42
Consistent123 [$*{}^{*}$ 17.13 0.7821 0.1255 1.55 18.02 20.11 0.8553 0.0716 3.24 20.08
Ours (Free3D) 18.23 0.8090 0.0996 1.34 8.67 21.13 0.8686 0.0619 2.85 9.10
Table 2: Generalizable results on unseen datasets, including OmniObject3D [81] and GSO [20], with 6,000 and 1,030 3D instances, respectively. Note that, although Zero123-XL [17] is trained on a larger dataset, and shows better generalizability, the proposed Free3D still significantly outperforms it with precise pose estimation for target views.
4.2 Assessing Quality
Quantitative comparison.
We first evaluate Free3D and other methods on the Objaverse dataset [18, 43], where the model is trained. Unlike previous work [43, 77, 44] which consider only a subset of the test data due to expensive post-processing, we use the entire test set of 7,729 3D objects. Quantitative results in Tab. 1 show that Free3D outperforms state-of-the-art models. This includes Zero123-XL [17] and SyncDreamer [44], which are trained on a much larger dataset and with explicit 3D volume representation, respectively. While the concurrent Consistent123 [77] also utilizes a form of multi-view diffusion with cross-view attention, our Free3D significantly improves the quantitative results ($16\%$ relative improvement on LPIPS) on the large evaluation dataset. This suggests that our RCN layer is the primary reason for the observed improvements.
Refer to caption(a) Input View(b) Zero-1-to-3 [43](c) Consistent123 [77](d) Ours Free3D
Figure 6: Qualitative new-view synthesis comparisons. Zero-1-to-3 [43] generates diverse details (e.g. various ears and tails) for different views in one sampling. Consistent123 [77] improves it through the multi-view diffusion with the cross-view attention. However, it still requires training additional NeRF for 3D reconstruction. For more comparisons, see the supplemental videos for better visualisation.
Method Objaverse [18] GSO [20]
[…]
$\mathbb{A}$ Baseline Zero-1-to-3 [43] 19.65 0.8462 0.0938 1.52 22.10 19.65 0.8501 0.0758 3.24 33.15
$\mathbb{B}$ + Input Ray Embeddings 20.21 0.8550 0.0858 1.26 16.63 20.49 0.8617 0.0677 3.01 22.08
$\mathbb{C}$ + Multi-Scale Ray Emb. 20.56 0.8609 0.0797 1.30 15.94 20.50 0.8615 0.0667 2.98 20.09
$\mathbb{D}$ + RCN 20.78 0.8620 0.0784 1.21 15.67 21.13 0.8686 0.0619 2.85 18.48
$\mathbb{E}$ + Pseudo-3D attention 20.81 0.8620 0.0781 1.25 14.76 21.20 0.8697 0.0617 2.86 17.39
$\mathbb{F}$ $\mathbb{E}$ + noise sharing 11.39 9.10
Table 3: Ablations of Free3D design choices. Here, to reduce the computational cost and testing time, we only evaluate 2,000 instances in Objaverse dataset, instead of running the whole dataset with 7,729 for 50-frame $360^{\circ}$ video rendering. Therefore, the PPLC score is slight different from the values reported in Tab. 1.
Qualitative comparison.
Qualitative results on various categories are visualized in Fig. 4. Free3D achieves better results even under challenging viewpoints with very different categories. SyncDreamer [44] uses an explicit volumetric representation representation, but can only generate views with fixed elevation ($30^{\circ}$), leading to worse results on synthesizing arbitrary target views. While the concurrent Consistent123 [77] aims to improve rendering consistency with a version of multi-view attention, they cannot directly improve pose accuracy. The Zero123-XL [17] learns to capture pose better than the baseline [43] by training on the larger Objaverse-XL, but the pose is still not very accurate. Because of the RCN layer, Free3D shows no such pose errors and results in better images.
4.3 Assessing Generalization
In Tab. 2 and Fig. 5, we validate the ability of Free3D to generalize to datasets not seen during training. This includes the OmniObject3D and the GSO datasets, with 6,000 scanned objects in 190 categories and 1,030 scanned objects in 17 categories, respectively. For this result, we directly test all trained models without any fine-tuning.
Quantitative comparison.
In Tab. 2, Free3D, which uses the same training set as Zero-1-to-3 [43], outperforms the baseline and all state-of-the-art variants, including Zero123-XL [17], which is trained on a larger 3D dataset, and SyncDreamer [44], which utilizes a heavier 3D volumetric representation. SyncDreamer improves the baseline on OmniObject3D, which has small elevation views change, while achieving worse results on GSO with random elevation angles. The concurrent Consistent123 [77] is also trained multiple views jointly, with additional cross-view attention. However, their relative improvement is limited. Table 2 shows that Free3D achieves very substantial improvements on all instantiations on both OmniObject3D and GSO datasets. This further indicates that ray conditioning can successfully improve the pose accuracy of the target view.
Qualitative comparison.
A qualitative comparison is given in Fig. 5 (more in supplemental Figs. C.3 and C.4). Though Free3D is only trained on objaverse dataset, it works quite well in the open-set setting. Moreover, it shows significantly better results than all state-of-the-art models, and even better than the Zero123-XL [17], a concurrent model trained on much larger datasets.
4.4 Assessing 3D consistency
To assess 3D consistency, we render a $360^{\circ}$ video with the fixed elevation angle $\theta:=0^{\circ}$ and $50$ azimuth angles uniformly sampled in $[0^{\circ},360^{\circ}]$. The quantitative results are reported in Tabs. 1 and 2 using the proposed PPLC score, and the qualitative results are shown in Figs. 1 and 6. Consistent123 [77] still requires training additional NeRF to achieve 3D consistent video, while the directly rendered videos are still flickering. SyncDreamer [44] obviously improves view consistency by using an explicit 3D volume representation. However, it is trained only on a fixed elevation angle and can generate only 16 fixed frames for a video in its code. Interestingly, Free3D, by using multi-view attention and multi-view noise sharing is nearly as effective (while being much cheaper). This is also partly due to ray conditioning, which can capture more precise poses of the target view and thus reduce ambiguity. The additional results are provided as videos in the supplemental materials.
4.5 Ablation Study
We run a number of ablations to analyse Free3D, testing also the variants described in Sec. 3. Results are shown in Tab. 3 and discussed in detail next.
Our baseline configuration (denoted $\mathbb{A}$) is the same as Zero-1-to-3 [43], which is derived from the SD model [57] by replacing text conditioning with source image and target pose conditioning. Then, in $\mathbb{B}$ we extend this model with ray conditioning by concatenating the ray embeddings to the source image $x^{\text{src}}$. This alone improves the performance dramatically in both Objaverse and GSO. In $\mathbb{C}$ we test injecting ray conditioning into each level of the diffusion UNet $\epsilon_{\theta}$, but the generalizable performance remains similar to $\mathbb{B}$ in the unseen GSO dataset. In $\mathbb{D}$, we test instead the RCN layer, which results in further improvements on not only Objaverse dataset, but also on the new GSO. In $\mathbb{E}$, we add multi-view attention to exchange information between different frames. As expected, this does not improve the metrics that measure the quality of individual views, but it slightly improves the PPLC score, measuring consistency. In $\mathbb{F}$, we further add multi-view noise sharing, which significantly enhances the consistency between different views. while preserving the quality of single-view rendering. Rendered videos are provided in the supplement.
5 Conclusion
We have introduced Free3D, an open-set single-view NVS method with state-of-the-art performance on various categories, yet bypass the requirement of building on heavy 3D representation or training additional auxiliary 3D models. It is a simple approach that (i) obtains data prior from an off-the-shelf pre-trained 2D image generator, (ii) injects ray conditioning utilizing the new RCN layer to accurately code for the target pose, and (iii) combines that with multi-view attention and noise sharing to improve multi-view consistency. Experimental results show that Free3D significantly outperforms recent and concurrent state-of-the-art NVS models without incurring the cost of utilizing a 3D representation. We hope that Free3D will serve as a new strong baseline for single-image NVS and inspire future research in this area.
Acknowledgements.
This research is supported by ERC-CoG UNION 101001212. Many thanks to Stanislaw Szymanowicz, Edgar Sucar, and Luke Melas-Kyriazi of VGG for insightful discussions and Ruining Li, Eldar Insafutdinov, and Yash Bhalgat of VGG for their helpful feedback. We would also like to thank the authors of Zero-1-to-3 [43] and Objaverse-XL [17] for their helpful discussions.
Ethics.
We use the Objaverse [18], OmniObject3D [81], and Google Scanned Object datasets (GSO) [20] following their terms and conditions. These datasets contain synthetic or scanned 3D objects, but, as far as we could determine, no personal data. For further details on ethics, data protection, and copyright please see https://www.robots.ox.ac.uk/~vedaldi/research/union/ethics.html.
Supplementary Material
The supplementary materials are organized as follows:
  • A video to illuminate our work and the rendered videos.
  • Introduction for the baseline diffusion model.
  • Experiment details.
  • Results for more single view NVS.
Appendix A Background: Diffusion Generators
In order to achieve sufficient generalization to operate in an open-set category setting, Free3D builds on a pre-trained 2D image generation, and specifically Stable Diffusion (SD) [57]. SD is a Latent Diffusion Model (LDM) trained on billions of text-image pairs from LAION-5B [62]. It consists of two stages. The first stage embeds the given image $x_{0}\in\mathbb{R}^{H\times W\times 3}$ in a latent space $z\in\mathbb{R}^{\frac{H}{f}\times\frac{H}{f}\times c}$ through an autoencoder $\mathcal{E}:x\mapsto z$, paired with a decoder $\mathcal{D}:z\mapsto x$, which reconstructs the image ($x=\mathcal{D}\circ\mathcal{E}(x)$). The second stage uses diffusion to model the distribution $p(z|y)$ over such latent codes, where $y$ lumps any conditioning information (e.g., text, image, or viewpoint). Diffusion involves a forward noising process that gradually perturbs the given latent $z_{0}=z$ by adding the Gaussian noise $\epsilon$ in a Markovian fashion:
$z_{t}=\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,\quad\epsilon\sim\mathcal{N}(0,I),$ (A.1)
producing a sequence $z_{t}$, $t=1,\dots,T$, and $\bar{\alpha}_{t}:=\prod_{s=1}^{t}\alpha_{s}$, $\alpha_{t}:=1-\beta_{t}$ denote the noise strength at different steps. $\{\beta_{t}\}_{t=1}^{T}$ is a pre-defined variance schedule. Ultimately, $p(z_{T}|y)$ is approximately normal; we can thus easily sample $z_{T}$, and then go back to $z_{0}$ via the backward denoising process using the predicted noise:
$z_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(z_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha_{t}}}}\epsilon_{\theta}(z_{t},t,y)\right)+\sigma_{t}\epsilon,$ (A.2)
where $\epsilon_{\theta}$ is typically an UNet [19], and $\{\sigma_{t}\}_{t=1}^{T}$ is another control of noise $\epsilon$, which is also a pre-defined schedule corresponding to the schedule $\beta_{t}$ and introduces uncertainty for the synthesis of different views. Similar to the vanilla DDPM [29], SD uses the following training objective to optimize the UNet $\epsilon_{\theta}$:
$\mathcal{L}=\mathbb{E}_{z_{0},y,\epsilon\sim\mathcal{N}(0,I),t}\left[\lVert\epsilon-\epsilon_{\theta}(z_{t},t,y)\rVert^{2}_{2}\right],$ (A.3)
Appendix B Experiment Details
The Stable Diffusion (SD), originally trained for text-to-images generation, requires adaptation to suit image-conditional NVS tasks. Following Zero-1-to-3 [43], we utilize the image-to-image Stable Diffusion checkpoints [^3] https://huggingface.co/spaces/lambdalabs/stable-diffusion-image-variations. Our baseline code is built upon the Zero-1-to-3 [43] [^4] https://github.com/cvlab-columbia/zero123. Hyperparameters are configured in accordance with the default settings of the baseline code. The ray conditioning normalisation (RCN) is incorporated into each ResNet block within the diffusion Unet $\epsilon_{\theta}$, while the pseudo-3D cross-attention is introduced after the original CLIP-conditional cross-attention layer (as illustrated in Fig. 2).
Instead of directly providing $\bm{r}_{uv}=(\bm{o}\times\bm{d}_{uv},\bm{d}_{uv})$ to the network for modulating the features, we embed them into higher-dimensional features, following the approach of NeRF [48] and LFN [66]. In particular, we employ the element-wise mapping $\bm{r}\mapsto[\bm{r},\sin(f_{1}\pi\bm{r}),\cos(f_{1}\pi\bm{r}),\cdots,\sin(f_{K}\pi\bm{r}),\cos(f_{K}\pi\bm{r})]$, where $K$ is the number of Fourier bands, and $f_{k}$ is equally spaced to the sampling rate. In all experiments, $K$ is set as 6, leading to $678=2\times 3\times K_{\bm{o}}+3+2\times 3\times K_{\bm{d}}+3=(6+6)\times 6+$ 6 dimensional features (as depicted in Fig. 2(a)).
Training Details.
Our model was trained on 4$\times$ A40 48GB GPUs in two stages: i) We first finetuned the model with RCN, utilizing a batch size of 256 for 3 days on random camera viewpoints, enhancing the pose accuracy for target views. ii) Subsequently, the pseudo-3D cross-attention was finetuned on the 4 nearest views, employing a batch size of 192 for 2 days. In the second stage, different views from one instance were perturbed by adding noise from the same time step $t$.
In an alternative approach during the first stage, we initially attempted to jointly train the pseudo-3D cross-attention with random camera viewpoints. However, the performance is worse than the configuration $\mathbb{D}$. We believe this is because the camera viewpoints have a large gap along these random views in the rendered datasets, making it harder to calculate the similarity across these frames. In all experiments, we use AdamW with a learning rate of $10^{-5}$ for the old parameters in the original diffusion Unet $\epsilon_{\theta}$ and a 10×10$\times$10 × larger learning rate for new parameters, namely the parameters for RCN and Pseudo-3D cross-attention.
Inference Details.
At the testing phase, we configure the diffusion model with a sampling step set to $T=50$. The computational time for rendering a novel view using our proposed Free3D is approximately 3 seconds, utilizing an A6000 GPU. For a fair comparison, all models are evaluated on the same A6000 GPU employing the same batch size of 4. This batch size is chosen due to the operational constraints of syndreamer [44], which can only run such a small size. Additionally, we also utilize the CFG with a scale $s=3$ to guide the rendering for each target view.
${}^{\circ}$ Video Rendering.
To render a ${}^{\circ}$ video, we establish a circle trajectory by uniformly subdividing the azimuth $\phi$ into discrete intervals of $\frac{2\pi}{50}=7.2^{\circ}$, while the elevation $\theta$ and the distance $z$ remain fixed. For each 3D instance, we replicate the same latent variable $z_{T}$ over 50 frames, which can minimize temporal flickering across different views. Additionally, we also set the parameter $\sigma_{t}$ in Eq. A.2 to zero, thereby further mitigating uncertainty introduced by varying noise patterns.
Appendix C More Visual Results
More results on Objaverse NVS.
In Figs. C.1 and C.2, we present more visual comparisons on Objaverse datasets [18] that given one input image and the target viewpoint, all models render the target novel view. This is an extension of Fig. 4 in the main paper.
Here, all examples shown come from the corresponding test-set following the split, as in Zero-1-to-3 [43]. These examples are good evidence that our Free3D is suitable for open-set categories NVS, where it can generate semantically reasonable content with visually realistic appearances across various categories. More importantly, compared to existing state-of-the-art methods, the Free3D provides better results with a more precise pose for the target novel view. This observation suggests that the RCN is able to provide better viewpoint perception for the NVS.
More results on OmniObject3D and GSO NVS.
In Figs. C.3 and C.4, we show additional comparison results on OmniObject3D [81] and GSO [20] datasets, respectively. This is an extension of Fig. 5 in the main paper, which demonstrates the generalizability of our Free3D on unseen datasets encompassing various categories.
As can be seen from these results, although the baseline Zero-1-to-3 [43] provides visually realistic appearances for all objects, the content is not always reasonable, and the pose is inaccurate in many cases. This indicates the global language token embedding with elevation $\theta$, azimuth $\phi$, and distance $z$ is not so precise for the network to interpret and utilize the camera viewpoints. While the Zero123-XL [17] and consistent123 [77] enhance the quality by training on a larger dataset and employing multi-view diffusion, respectively, they do not directly deal with the camera pose perception. In contrast, our Free3D leverages the per-pixel ray conditioning as well as the modulating, which significantly improves the pose perception accuracy.
Refer to caption(a) Input View(b) Target View(c) Zero123[43](d) Zero123-XL[17](e) SyncDreamer[44](f) Consistent123[77](g) Ours Free3D
Figure C.1: Qualitative comparisons on Objaverse dataset. Given the exact target pose, the proposed Free3D significantly improves the pose precision compared to existing state-of-the-art methods.
Refer to caption(a) Input View(b) Target View(c) Zero123[43](d) Zero123-XL[17](e) SyncDreamer[44](f) Consistent123[77](g) Ours Free3D
Figure C.2: Qualitative comparisons on Objaverse dataset. Given the exact target pose, the proposed Free3D significantly improves the pose precision compared to existing state-of-the-art methods.
Refer to caption(a) Input View(b) Target View(c) Zero123[43](d) Zero123-XL[17](e) SyncDreamer[44](f) Consistent123[77](g) Ours Free3D
Figure C.3: Qualitative comparisons on OminiObject3D dataset. Given the exact target pose, the proposed Free3D significantly improves the pose precision compared to existing state-of-the-art methods.
Refer to caption(a) Input View(b) Target View(c) Zero123[43](d) Zero123-XL[17](e) SyncDreamer[44](f) Consistent123[77](g) Ours Free3D
Figure C.4: Qualitative comparisons on GSO dataset. Given the exact target pose, the proposed Free3D significantly improves the pose precision compared to existing state-of-the-art methods.